a wire of length 62.0cm and mass 130.0g is suspended by a pair of flexible leads in a magnetic field of 440mt. find the magnitude and direction of the current in the wire required to remove the tension in the supporting leads
1
Expert's answer
2020-10-12T07:50:25-0400

Given parameters,

length l=62cm=0.62m

mass m=130g=0.13kg

magnetic field B=440mt=0.44t

Force exerted by gravity on the wire=mg

="0.13\times 9.8"

=1.274 N


To remove the tension in a wire ,force mg must be balanced by magnetic force

"F_b=mg"

"Bil=mg" ( where i is current in the wire)

"0.44\times i\times 0.62=1.274"


"i \times 0.2728=1.274"


"i=\frac{1.274}{0.2728}"


i =4.67A (from left to right in the wire)